# Revisiting Reweighted Risk for Calibration: AURC, Focal Loss, and Inverse Focal Loss

### v1

@@ 3.1 Weighted Risk Functions
$\mathcal{R}(f)=\mathbb{E}[\ell(\boldsymbol{p},\boldsymbol{y})],$ (1)
@@
Let $\boldsymbol{x}$ be an input vector, $\gamma\geq 0$ a focusing parameter, and $f$ a model mapping from the feature space into the probability simplex. The focal loss is defined as:
@@
Let $\boldsymbol{x}$ be an input vector, $\gamma\geq 0$, and $f$ a model mapping into the probability simplex. The inverse focal loss is defined as:
@@ 3.2 Selective classification
Let $\boldsymbol{p}=f(\boldsymbol{x})$ be the predicted probability vector produced by the model. The cumulative distribution function (CDF) of the CSF $g(\boldsymbol{p})$ is defined as:
@@
This formulation reveals that AURC can be interpreted as a reweighted risk, where each sample is weighted by $w(\boldsymbol{p})=-\ln(1-G(\boldsymbol{p}))$, a function that depends on the CSF. Common choices for the CSFs include the Maximum Softmax Probability (MSP), i.e., $g(\boldsymbol{p})=\max_{i=1}^{k}p_{i}$, or the entropy of the logits prior to the softmax layer.
@@ 3.3 Calibration Errors
A model $f:\mathcal{X}\rightarrow\Delta^{k}$ is calibrated if and only if $f(\boldsymbol{x})=\mathbb{E}[\boldsymbol{y}\mid f(\boldsymbol{x})]$.
@@
where the expectation is taken with respect to $\mathbb{P}(\boldsymbol{x},\boldsymbol{y})$.
When $\rho=1$, this quantity is referred to as the Expected Calibration Error (ECE) [17]. In practice, the true value of $\text{CE}_{\rho}(\cdot)$ is unobservable, as it depends on the unknown joint distribution $\mathbb{P}(\boldsymbol{x},\boldsymbol{y})$. To address this, Guo et al. [17] proposed an empirical estimator based on a finite dataset. This estimator employs a binning scheme $B$ to approximate the true difference between the predicted confidence and the conditional probability $\mathbb{E}[\boldsymbol{y}\mid f(\boldsymbol{x})]$ by computing empirical averages within each bin. We denote this estimator as $\widehat{\text{ECE}}_{B}$.
@@
To avoid potential confusion, we emphasize here the empirical estimator of ECE measures the top-label calibration error instead of the calibration error w.r.t. a fixed class. The class-wise calibration error (cwECE) [30] measures the expected calibration errors across all classes.
@@
$\text{cwECE}_{\rho}(f)=\frac{1}{k}\Big{(}\sum_{c=1}^{k}\mathbb{E}\left[\left|f_{c}(\boldsymbol{x})-\mathbb{P}(y^{\prime}=c\mid f_{c}(\boldsymbol{x}))\right|^{\rho}\right]\Big{)}^{1/\rho},$ (10)
@@ 4 A Differentiable Loss Function Derived from Selective Classification
In this section, we begin by examining lower bounds for two widely used calibration metrics—ECE and class-wise ECE—as presented in Prop. 4.14.3. These results provide theoretical insight into how calibration error can be reduced for a pre-trained model and motivate a deeper exploration of the connection between selective classification and calibration. Building on these insights, we propose a differentiable loss function derived from the AURC metric in selective classification. The formulation is fully differentiable—enabled by the SoftRank method—and supports arbitrary CSFs.
@@ 4.1 Linking Calibration Error Minimization to Selective Classification
Let $f:\mathcal{X}\!\to\!\Delta^{k}$ be the classifier and $\widehat{\operatorname{err}}(f)$ be its empirical error rate on a finite set $\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$ given by $\widehat{\operatorname{err}}(f)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\bigl{(}\hat{y}_{i}\neq y_{i}^{\prime}\bigr{)}$ where $\hat{y}_{i}=\arg\max\limits_{c\in[k]}f_{c}(x_{i})$. For the top-label ECE computed over all binning schemes $B\in\mathcal{B}$,
@@
Let $g$ be a CSF that perfectly separates correct and incorrect predictions via threshold $\tau$, and let $f:\mathcal{X}\rightarrow\Delta^{k}$. Then a modified model $f^{\prime}$ can be constructed such that:
$\sup_{B\in\mathcal{B}}\widehat{\operatorname{cwECE}}_{B}(f^{\prime})\leq\sup_{B\in\mathcal{B}}\widehat{\operatorname{cwECE}}_{B}(f),$ (16)
with:
@@
Let $f:\mathcal{X}\to\Delta^{k}$ be a classifier, and let $1\leq\rho\leq 2$. Then:
@@ 4.2 Regularized AURC Loss Function
Given the preceding results, reducing calibration error requires lowering the maximum predicted probability for low-confidence samples—a strategy that echoes the "abstain" mechanism central to selective classification. Since the AURC metric originates from the selective classification framework and employs a weighting scheme similar to that of inverse focal loss (see Fig. 1), such reweighting schemes are theoretically well-suited for calibration error minimization. Motivated by this, we propose the regularized AURC (r-AURC) loss to examine whether this reweighting approach can effectively enhance model calibration. An asymptotically unbiased Monte Carlo estimator [46] for the population AURC defined in Eq. (7) is given by:
@@
This loss function typically relies on sorting based on the confidence score function (CSF), incurring a computational cost of $\mathcal{O}(n\log(n))$ computation cost, which is generally more efficient than kernel-based calibration error losses that require quadratic time. To better understand the effect of CSF on AURC metric, we define the perfectly aligned CSF.
@@
To this end, we consider the following constrained optimization framework to learn a model $f$ that is both accurate and well-calibrated: $f=\arg\min_{f\in\mathcal{F}}\mathcal{R}(f),\quad\text{subject to}\quad\text{AURC}_{g}(f)\leq B,$ where $\mathcal{R}(f)$ denotes the expected predictive risk, and $B>0$ is a user-defined budget. The corresponding Lagrangian relaxation of this constrained objective is given by:
$f=\arg\min_{f\in\mathcal{F}}\left((1-\lambda)\,\mathcal{R}(f)+\lambda\,\text{AURC}_{g}(f)\right),$ (22)

### v2

@@ 3.1 Weighted Risk Functions
$\mathcal{R}(f)=\mathbb{E}[\ell(f\left(\boldsymbol{x}\right),\boldsymbol{y})],$ (1)
@@
Let $\boldsymbol{p}$ be a predicted probability vector, and $\gamma\geq 0$ a focusing parameter. The focal loss is defined as:
@@
Let $\boldsymbol{p}$ be a predicted probability vector, $\gamma\geq 0$, and $f$ a model mapping into the probability simplex. The inverse focal loss is defined as:
@@ 3.2 Selective classification
Let $\boldsymbol{p}=f(\boldsymbol{x})$ be a predicted probability vector produced by the model $f$ with fixed input $\boldsymbol{x}$. Denote with $\boldsymbol{p}^{\prime}=f\left(\boldsymbol{x}^{\prime}\right)$ the output random variable with input random variable $\boldsymbol{x}^{\prime}$. The cumulative distribution function (CDF) of the CSF $g(\boldsymbol{p})$ is defined as:
@@
This formulation reveals that AURC can be interpreted as a reweighted risk, where each sample is weighted by $w(\boldsymbol{p})=-\ln(1-G(\boldsymbol{p}))$, a function that depends on the CSF. Common choices for the CSFs include the Maximum Softmax Probability (MSP), i.e., $g(\boldsymbol{p})=\max_{i=1}^{k}p_{i}$, or the negative entropy of the logits prior to the softmax layer as described in Table 3.
@@ 3.3 Calibration Errors
A model $f:\mathcal{X}\rightarrow\Delta^{k}$ is calibrated if and only if $f(\boldsymbol{x})=\mathbb{E}[\boldsymbol{y}\mid f(\boldsymbol{x})]$ almost surely based on the input random variable $\mathbf{x}$.
@@
where the expectation is taken with respect to $\boldsymbol{x}$.
When $\rho=1$ and for binarized outcomes, this quantity is referred to as the Expected Calibration Error (ECE) [38, 17], i.e.,
$\operatorname{ECE}\left(f\right)=\mathbb{E}\left[\lvert\max_{c}f_{c}(\boldsymbol{x})-\mathbb{P}(\boldsymbol{y}=\arg\max_{c}f_{c}(\boldsymbol{x})\mid\max_{c}f_{c}(\boldsymbol{x}))\rvert\right].$ (9)
In practice, the true value of $\text{CE}_{\rho}(\cdot)$ is unobservable, as it depends on the unknown conditional distribution $\mathbb{P}(\boldsymbol{y}\mid f\left(\boldsymbol{x}\right))$. To address this, Naeini et al. [38] proposed an empirical estimator based on a finite dataset. This estimator employs a binning scheme $B$ to approximate the true difference between the predicted confidence and the conditional probability $\mathbb{E}[\boldsymbol{y}\mid f(\boldsymbol{x})]$ by computing empirical averages within each bin. We denote this estimator as $\widehat{\text{ECE}}_{B}$.
@@
To avoid potential confusion, we emphasize here the empirical estimator of ECE measures the top-label calibration error instead of the calibration error w.r.t. a fixed class. The class-wise calibration error (cwECE) [30, 29] measures the expected calibration errors across all classes.
@@
$\text{cwECE}_{\rho}(f)=\Big{(}\frac{1}{k}\sum_{c=1}^{k}\mathbb{E}\left[\left|f_{c}(\boldsymbol{x})-\mathbb{P}(y^{\prime}=c\mid f_{c}(\boldsymbol{x}))\right|^{\rho}\right]\Big{)}^{1/\rho},$ (11)
@@ 4 A Differentiable Loss Function Derived from Selective Classification
In this section, we begin by examining lower bounds for two widely used calibration metrics—ECE and class-wise ECE—as presented in Prop. 4.14.3. These results provide theoretical insight into how the calibration error can be reduced for a pre-trained model and motivate a deeper exploration of the connection between selective classification and calibration. Building on these insights, we propose a differentiable loss function derived from the AURC metric in selective classification. The formulation is fully differentiable—enabled by the SoftRank method—and supports arbitrary CSFs.
@@ 4.1 Linking Calibration Error Minimization to Selective Classification
Let $f:\mathcal{X}\!\to\!\Delta^{k}$ be the classifier and $\widehat{\operatorname{err}}(f)$ be its empirical error rate on a finite set $\{(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\}_{i=1}^{n}$ given by $\widehat{\operatorname{err}}(f)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\bigl{(}\hat{y}_{i}\neq y_{i}^{\prime}\bigr{)}$ where $\hat{y}_{i}=\arg\max\limits_{c\in[k]}f_{c}(x_{i})$. For the top-label ECE computed over all binning schemes $B\in\mathcal{B}$ holds that
@@
Let $g$ be a CSF that perfectly separates correct and incorrect predictions via the threshold $\tau$, and let $f:\mathcal{X}\rightarrow\Delta^{k}$. Then, a modified model $f^{\prime}$ can be constructed such that
$\sup_{B\in\mathcal{B}}\widehat{\operatorname{cwECE}}_{B}(f^{\prime})\leq\sup_{B\in\mathcal{B}}\widehat{\operatorname{cwECE}}_{B}(f)$ (17)
with
@@
Let $f:\mathcal{X}\to\Delta^{k}$ be a classifier, and let $1\leq\rho\leq 2$. Then,
@@ 4.2 Regularized AURC Loss Function
Given the preceding results, reducing calibration error requires lowering the maximum predicted probability for low-confidence samples—a strategy that echoes the "abstain" mechanism central to selective classification. Since the AURC metric originates from the selective classification framework and employs a weighting scheme similar to that of inverse focal loss (see Fig. 1), such reweighting schemes are theoretically well-suited for calibration error minimization. Motivated by this, we propose the regularized AURC (r-AURC) loss to examine whether this reweighting approach can effectively enhance model calibration. An asymptotically unbiased Monte Carlo estimator [47] for the population AURC defined in Eq. (7) is given by
@@
This loss function typically relies on sorting based on the confidence score function (CSF), incurring a computational cost of $\mathcal{O}(n\log(n))$, which is generally more efficient than kernel-based calibration error losses that require quadratic time. To better understand the effect of CSF on AURC metric, we define the perfectly aligned CSF.
@@
To this end, we consider the following constrained optimization framework to learn a model $f$ that is both accurate and well-calibrated: $f^{*}=\arg\min_{f\in\mathcal{F}}\mathcal{R}(f),\quad\text{subject to}\quad\text{AURC}_{g}(f)\leq B,$ where $\mathcal{R}(f)$ denotes the expected predictive risk, and $B>0$ is a user-defined budget. The corresponding Lagrangian relaxation of this constrained objective is given by:
$f^{*}=\arg\min_{f\in\mathcal{F}}\left((1-\lambda)\,\mathcal{R}(f)+\lambda\,\text{AURC}_{g}(f)\right),$ (23)